# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v2

@@ Abstract
In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. Our method provides marked improvements in the constants, especially with regards to the latter. These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s. Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
@@ 1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [26], [4], [9], [25, (3.14)] for more information on the topic.
Trudgian in [26] gives an explicit version of results in [25, §3] originally due to Landau, obtaining bounds of order $O(\log t)$. This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$. However, such estimates tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
The method in [25, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$. It relies on two applications of the Borel–Carathédory theorem. The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$. As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [13, Lemma 3.2].
While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [26], we take the opportunity to point out an error in that paper. In the quantity $A$ of [26, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. In this article, we have recovered the erroneous bounds and even slightly improved upon them.
@@ 2.1 Properties of the Riemann zeta function
We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$. Similarly we denote the Riemann hypothesis by $RH$, by which we mean all zeros $\rho$ have $\beta=1/2$.
@@
where $Y(W_{0},10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
@@
where $Y^{\prime}(W_{0},10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
@@ 3 Some preliminaries
Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2]. The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [24, §3.71]).
@@
Combining this with the bound on $J$ and (3) proves the theorem. ∎
@@
Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [27, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
@@
[28, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
@@
The proof is analogous to that of [27, Lemma 2.1], except that we are more precise in estimating (7) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
Note that different to the proof of [27, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (8)) in $\sigma$. Additionally, we found that further splitting the interval complicates (8) and raises the threshold of $t$ for the lemma to be valid. ∎
@@
We now refer to Lemma 4, taking the maximum of (7) at $\sigma=1/2$. It has been shown in [29, p. 15] that $Q_{0}=1.31$ is permissible in (6) together with
@@
We use same argument in the first paragraph of [28, p. 2], which modifies Ford’s proof in [10]. The only difference is that we now apply the argument to [2, §5], which itself is an improvement on Ford’s result. This proves the lemma. ∎
@@
[22, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
@@
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$. Note that (12) holds for $\sigma>0$ by analytic continuation (see [25, Theorem 3.5]).
@@
We thus have
$|\zeta(s)|$ $\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}+\frac{1}{64N^{4}}+\frac{1}{t}+\frac{t^{2}+\sigma_{1}^{2}+2(t+\sigma_{1})-1}{12N^{2}}\right).$
@@
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [27, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [15]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
[21, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
@@
It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [21, Lemma 5.3]. Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (19) are available, but are more complicated to work with.
@@
Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [25, § 3.3] and [18]). That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
@@ 4.1 Assuming a partial Riemann hypothesis
$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$ (31)
@@ 4.2 Above the Riemann height
Let $d$ and $W\geq W_{0}$ be positive parameters, subject to (35), (36), and (44). Let also $10<\alpha<$ 1 be a parameter defined in (44), and depending on $d$ and $W$. In a region
@@ 4.3 Combining both cases $t\leq H$ and $t\geq H$
5 Bounds on $|1/\zeta(s)|$
Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [25, Theorem 3.11]). It revolves around estimating the right-hand side of
@@ 5.2 Above the Riemann height
Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log H\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W\geq W_{0}$ is chosen such that $d$ satisfies (36). Then in the region
@@ 5.3 Combining both cases $t\leq H$ and $t\geq H$
$t_{0}$ $Y(W_{0},t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
$10$ $8.101$ $0.307574$ $1.133578$ $3.071916$
$10^{2}$ $4.339$ $0.542834$ $1.388735$ $3.417735$
$10^{3}$ $3.632$ $0.648794$ $1.470345$ $3.459386$
$10^{4}$ $3.264$ $0.725027$ $1.834773$ $3.463610$
$10^{5}$ $3.021$ $0.787051$ $1.447903$ $3.464020$
$10^{7}$ $2.711$ $0.885013$ $1.30488$ $3.464071$
$10^{9}$ $2.518$ $0.960532$ $1.322282$ $3.464052$
$10^{11}$ $2.383$ $1.021393$ $1.511713$ $3.464182$
Table 5: Values for $Y(W_{0},t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}$ and $W=W_{0}$.
@@ 5.4 An asymptotic improvement
$\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
@@
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
@@
Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W\geq W_{0}$ is chosen such that $d$ satisfies (36). Then in the region
@@ 6 Concluding Remarks
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. As an example, for $W=W_{0}$, the condition (36) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$. The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us. In fact we only need a ‘weak’ version of $RH$, for instance in (24), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [25, §14], [23], [22], and [6] for more details).

### v3

@@ Abstract
In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. Our method provides marked improvements in the constants, especially with regards to the latter. These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s. Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement within the classical zero-free region: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
@@ 1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [27], [4], [9], [26, (3.14)] for more information on the topic.
Our main results will be presented in detail in Section 2, but here we highlight two of the more significant ones. In particular, Corollaries 4 and 5 of this article show respectively that for
$\sigma\geq 1-\frac{1}{5.5587\log t}\qquad\text{and}\qquad t\geq 1000,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 3.632\log t,\qquad\text{and}\qquad\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 9.842(\log t)^{11/12}.$
Now we give a brief history of results on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$. Trudgian in [27] gives an explicit version of results in [26, §3] originally due to Landau, obtaining bounds of order $O(\log t)$. This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$. However, such estimates tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
The method in [26, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$. It relies on two applications of the Borel–Carathédory theorem. The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$. As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [13, Lemma 3.2].
While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [27], we take the opportunity to point out an error in that paper. In the quantity $A$ of [27, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. In this article, we have recovered the erroneous bounds and even slightly improved upon them.
@@
Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us get an asymptotic improvement in the case of $1/\zeta(s)$. In particular, working within zero-free regions of the forms
$\sigma\geq 1-\frac{1}{Z_{1}\log t},\qquad\sigma\geq 1-\frac{\log\log t}{Z_{2}\log t},\qquad\sigma\geq 1-\frac{1}{Z_{3}(\log t)^{2/3}(\log\log t)^{1/3}},$
for some positive constants $Z_{1},\,Z_{2},\,Z_{3}$, one usually proves upper bounds for $1/\zeta(s)$ of the order
$\log t,\qquad\frac{\log t}{\log\log t},\qquad(\log t)^{2/3}(\log\log t)^{1/3},$ (1)
respectively. Without making any changes to the zero-free region used, our new method would allow one to refine the results in (1), to
$(\log t)^{11/12},\qquad\frac{(\log t)^{11/12}}{(\log\log t)^{3/4}},\qquad(\log t)^{2/3}(\log\log t)^{1/4},$ (2)
respectively. As far as we are aware, results of these types were not previously known. In this paper, we prove an explicit version of the first bound in (2). An explicit version of the third bound in (2) is given in [18].
@@ 2.1 Properties of the Riemann zeta function
We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all non-trivial zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$. Similarly we denote the Riemann hypothesis by $RH$, by which we mean all non-trivial zeros $\rho$ have $\beta=1/2$.
@@
where $Y(5.5587,10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
@@
where $Y^{\prime}(5.5587,10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
@@ 3 Some preliminaries
Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2]. The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [25, §3.71]).
@@
Combining this with the bound on $J$ and (5) proves the lemma. ∎
@@
Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [28, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
@@
[29, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
@@
The proof is analogous to that of [28, Lemma 2.1], except that we are more precise in estimating (9) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
Note that different to the proof of [28, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (10)) in $\sigma$. Additionally, we found that further splitting the interval complicates (10) and raises the threshold of $t$ for the lemma to be valid. ∎
@@
We now refer to Lemma 4, taking the maximum of (9) at $\sigma=1/2$. It has been shown in [30, p. 15] that $Q_{0}=1.31$ is permissible in (8) together with
@@
We use same argument in the first paragraph of [29, p. 2], which modifies Ford’s proof in [10]. The only difference is that we now apply the argument to [2, §5], which itself is an improvement on Ford’s result. This proves the lemma. ∎
@@
[23, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
@@
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$. Note that (14) holds for $\sigma>0$ by analytic continuation (see [26, Theorem 3.5]).
@@
We thus have that $|\zeta(s)|$ is upper bounded by
$\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}+\frac{1}{64N^{4}}+\frac{1}{t}+\frac{t^{2}+\sigma_{1}^{2}+2(t+\sigma_{1})-1}{12N^{2}}\right).$
@@
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [28, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [15]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
[22, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
@@
It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [22, Lemma 5.3]. Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (21) are available, but are more complicated to work with.
@@
Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [26, § 3.3] and [19]). That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
@@ 4.1 Assuming a partial Riemann hypothesis
$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$
@@ 4.2 Above the Riemann height
Let $d$ and $W>W_{0}$ be positive parameters, subject to (36), (37), and (45). Let also $10<\alpha<$ 1 be a parameter defined in (45), and depending on $d$ and $W$. In a region
@@ 4.3 Combining both cases $t\leq H$ and $t\geq H$
5 Estimates for $|1/\zeta(s)|$
Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [26, Theorem 3.11]). It revolves around estimating the right-hand side of
@@ 5.2 Above the Riemann height
Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log H\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37). Then in the region
@@ 5.3 Combining both cases $t\leq H$ and $t\geq H$
$t_{0}$ $Y(5.5587,t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
$10$ $8.101$ $0.307606$ $1.330692$ $3.068541$
$10^{2}$ $4.339$ $0.542825$ $1.851563$ $3.417699$
$10^{3}$ $3.632$ $0.648814$ $1.282661$ $3.465621$
$10^{4}$ $3.264$ $0.725026$ $1.748467$ $3.463629$
$10^{5}$ $3.021$ $0.787049$ $1.208003$ $3.464054$
$10^{7}$ $2.711$ $0.885011$ $1.101137$ $3.464109$
$10^{9}$ $2.518$ $0.960512$ $1.272668$ $3.445896$
$10^{11}$ $2.383$ $1.021396$ $1.655004$ $3.464097$
Table 5: Values for $Y(5.5587,t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}$ and $W=5.5587$.
@@ 5.4 An asymptotic improvement
$\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
@@
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
@@
Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37). Then in the region
@@ 6 Concluding Remarks
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. As an example, for $W=5.5587$, the condition (37) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$. The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us. In fact we only need a ‘weak’ version of $RH$, for instance in (26), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [26, §14], [24], [23], and [6] for more details).